A 140 g baseball traveling at 40 m/s is hit with a
Answer #1

Using Newton's second law we can estimate the average force acting on the ball during the hit:

F = dP/dt = (mvf - mvi) / t

3400 = 0.14(vf - 40) / 0.004

vf = 380 m/s

b) F = ma

a = 3400 / 0.14 = 24285 m/s^2